There are two consumers on the market: Jim and Donna. Jim’s utility function is U(x, y) = xy. Donna’s utility function is U(x, y) = x 2 y. Income of Jim is mJ = 100 and income of Donna is mD = 150.

a) Find optimal baskets of Jim and Donna when price of y is Py = 1 and price of x is P. (PREVIOUSLY POSTED)

b) On separate graphs plot Jim’s and Donna’s demand schedule for x for all values of P. (PREVIOUSLY POSTED)

c) Compute and plot aggregate demand when Jim and Donna are the only consumers.
Answer #1

C.

P AD
2 25+50=75
5 10+20=30
10 5+10=15

Graph caption 8.4 2 3.6 2 X axis variable

Y axis represents price